OpenStudy (anonymous):

how could i find the Maclurin series id ln(1+5x). Please and thank you

OpenStudy (anonymous):

fap fap fap! :DD

Nnesha (nnesha):

now delete that @stephaniesosa52 sorry :(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok, i'll be half serious. Take derivative use the fact that 1/(1-x) = 1 + x + x^2 = ... then integrate. Done! fap fap fap! again :D litterally